# The unified non-disciplinary model of the operation of neurons

### v3

@@ Abstract
The paper introduces the unified, non-disciplinary theory of neuronal operation that explains all known experimental observations without contradictions. The competing electrical and thermodynamic disciplinary theories mutually cannot explain some observations of neuronal operation, simply because some quantities fall outside of the scope of the given discipline. Although the first principles of science are identical in the inanimate and life sciences, the abstractions and approximations that underlie concrete laws differ, so the actual laws depend on the environment. We challenge the universality of protein mechanisms for describing neuronal operation and show that, by using the appropriate approximations, the non-disciplinary cooperation of thermodynamics and electricity perfectly describes processes in living matter. We show that biology works on the boundary of the micro and macro views of physics, and in some cases, those views must be bridged. Most importantly, we explicitly consider the discrete nature of biological current, that the charge carriers are slow and they repel each other. That feature, overlooked in both mentioned theories, enables us to interpret the electrical and thermodynamic behavior of important biological objects and processes. The speed of biological currents changes by several orders of magnitude. It keeps the electrolyte in continuous movement, making some small regions of the volume responsible for the observed action potential. The cross-disciplinary model enables us to derive the correct electrical equations and thermodynamic description of the action potential. In a way demonstrating the need for the non-disciplinary discussion, we derive the thermodynamic Carnot-cycle of AP from measuring electrical parameters, calculate its energy consumption and efficiency, interpret the neuron’s partially reversible operation and the long-sought solution to the heat emission/absorption issue. We defy the claim that science cannot describe life: the underlying physics constrains biological phenomena, including cognitive ones.
@@ 1 Introduction
The "great journey into the unknown [understanding the brain]" [2] must begin at a much lower level: revisiting the fundamental phenomena, disciplines, laws, interactions, abstractions, omissions, and testing methods of science. Research must build on classical science, but be reinterpreted for living matter. There is no independent ’life science’, there is only science. It is based on the same ’first principles’ but using different abstractions and approximations for living and non-living matters, and having the appropriate relations between them. "There is a clear need for a tighter and more carefully managed integration and realignment of the work" [3]. Without aligning the knowledge elements along the first principles, the "integration between data and knowledge from different disciplines", lacks integration. However, even the review of the targeted Human Brain Project (HBP) summarized that "HBP is not developing with the expected level of integration and the project controls in place are not adequate to achieve this aim." [3]. We arrived at the boundaries of classical science disciplines and are moving now through terra incognita.
@@
Our model is undoubtedly a physical model, in the above sense. We provide a holistic, multi-level view of neuronal operation and describe (at least semi-)quantitatively the actual physical processes; this way, we achieve the promised ’new understanding’ (and the need for rewriting books in neuroscience). Although the classical neuron model aims to be an electrical model, it violates fundamental laws of electricity and conservation laws; furthermore, it cannot interpret phenomena from other disciplines (thermodynamics, mechanics, optics). The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We prove that no compromise is needed: the pressure and the voltage measurements observe the same physical process by different disciplinary means. Neuroscience forgot that neural currents consist of ions, which repel each other in a confined volume that contains an object with static charge; furthermore, that ions are charged and have mass, both of which are abstractions from different classical disciplines.
@@
Research already provided the detailed description of a correct ‘net electrical’ neuron model [5], the laws of motion for describing biology [6], and now provides a correct non-disciplinary model (promising a unified and scientifically complete electrical/thermodynamic model). The deeper reason for the apparent conflict between the two competing theoretical descriptions is science’s disciplinarity, as we discuss it separately in Appendix A, since it is somewhat philosophical and touches not only the fundamental concepts of physics but also the way of scientific thinking. Shortly, the disciplinary separation proved to be successful for most phenomena of inanimate science. Electrolytes are an exception. Thermodynamics (due to excluding a long-range interaction between its particles), excludes electric charge (and related quantities) from the quantities it studies. Electricity (due to implying ’electron cloud’ as a charge transmission method) similarly excluded that long-range interaction may exist between its discrete charge carriers; that is, that they may have thermodynamic features based on the Vlasov equations. This way, electricity excludes mass (and related quantities, such as heat) from the quantities it studies. Neither of the two disciplines, alone, can completely describe electrolytes (and so: biological systems): some of the system’s quantities remain outside the scope of the given discipline. The measurements are not complete: some quantities are not measured, so ’miracles’ can occur. The force analysis does not show a driving force, so some "new force or whatnot" [7] must be assumed (in physiology, "protein mechanisms"). Our non-disciplinary approach, however, can successfully describe the operation of a neuron, without a miracle.
@@
The structure of the paper is as follows. Section 2 presents the state-of-the-art of theoretical neuron modeling. A strongly simplified discussion in section 3, without mathematical, physical, and electrical details, presents the unified model. Electricity of ions and its application to the case of neurons is the subject of section 4, which includes the specifics of ion currents, the electrical features and operation of biological objects, and, first of all, the thermodynamics applied to neurons. Section 5 presents the physical mechanisms that, instead of a magic protein mechanism, define and maintain the membrane’s potential in resting and transient states. The importance of the action potential deserved a separate section 6, where the physical processes, their correct "net electrical" model, with the related mathematics, are described. The discussion culminates in section 7, which on one side derives the thermodynamic parameters describing producing Action Potential (AP) and their long sought calculation method, on the other provides evidence that only the cooperation of the competing disciplines can succeed: the thermodynamic parameters can be calculated by measuring parameters of electrical processes. The numerical estimations, serving as the base for some statements of the paper, are detailed in Appendix B.
@@ 2 State-of-the-art
Today, two major branches of disciplinary theories compete for being applied to describe the neuronal operation [8, 9, 10, 11, 12, 13]. Those cited references also compare those theories. The electrical view essentially stems from the (Nobel Prize-winning) work of Hodgkin and Huxley [14]. They could not make a perfect job, mainly due to the lack of discoveries made several decades later, including ours, about handling the finite speed of the biological ionic currents (and, due to that, the dynamic features), which drastically change their conclusions. In contrast with their empirical description, which delivered mathematically formulated measured observations, our discussion, although it follows essentially the same principles, reinterprets and pinpoints the used fundamental terms, sets up a physical model, and explains the physically underpinned processes. “However, the theory could not explain the physical phenomena such as reversible heat changes, density changes, and geometrical changes observed in the experiments” [11]. The thermodynamic view roots in the (possibly Nobel-prize-winning) idea of Heimburg and Jackson [15]. They proposed that the action potential is essentially a sound wave (a soliton). “However, there are several other questions that this has to answer like ion flow involvement in nerve signal propagation as stated by the Hodgkin and Huxley (HH) model and also the faster propagation in myelinated nerves than in unmyelinated” [11]. If we consider that the ion flow means charge propagation in a membrane tube where the specific capacity depends on the thickness of the wall (the myelin sheath) [16], the faster propagation is not mysterious anymore. Given that our (non-disciplinary) model natively connects charge and mass of ions [6], and it can explain the missing phenomena.
@@
The excellent textbook [17] separates the neuron membrane’s states into resting and transient states. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; we tell the second half in section 4. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential" [17], page 126. As we show in section 5, two different physical mechanisms operate in the resting and the transient states. A primary reason for misguiding neurophysiology was projecting the mechanism of the resting state to the transient state.
@@
The book [17] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, discussing only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization. We demonstrate in section 5 that classical methods of electricity enable us to calculate the membrane’s potential resulting from charge separation and polarization. Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [18]. However, we add that mainly physical processes and speed gradients control the processes. More precisely, inseparable thermodynamic and electrical processes keep the potential set by the inseparable electrical and thermodynamic backbone defined by the lipid/protein structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [19].
@@ 3 Operation in a nutshell
We have E. Schrödinger’s comments in mind when constructing the physical picture of neurons, that we must be cautious. We must also include the discoveries from the past seven decades, among others, that in the transient state, a neuron shall be modeled as a serial (i.e., differentiator-type) electrical oscillator, which has gated current inputs (instead of a parallel oscillator without gating inputs, which is a good model only in the resting state). We explicitly add that the charge carriers, the neuron works with, are ions (instead of electrons), and that the signal propagation mechanism is bioelectric (instead purely an electric field or mechanical wave propagation or "protein mechanisms"), in many cases even without having free carriers in the respective volume. (We do not consider the ideas that the associated mechanical, optical, etc. changes cause electrical phenomena. They reverse the true causality: they do not have a triggering cause; the view inherited from the narrative biology.) Electrochemical operation, especially that of the semipermeable membranes, requires special attention: charge carriers are "created" inside the biological matter. Omitting the precise physical background results in performing wrong measurements and the introduction of wrong concepts into biology.
@@
In the non-disciplinary physical model (the ’abstract physical neuron’, as we call it), a neuron (highly simplified) is a sphere with two lipid layers (membrane), a similarly insulating output tube (axon), and an incompressible, partially ionized electrolyte fluid with different composition and concentration inside and outside. They contain ions, simple chemical molecules, and complex biological formations. These components may have gradients and move at continuously changing speeds that can vary by several orders of magnitude during operation. The lipid layer is partially permeable (controlled or uncontrolled) for the components above (channels, mainly ion channels).
@@
The membrane represents a confined space in which ions behave differently from in an infinite space; furthermore, it significantly alters the properties of the few-nanometer-thick layer near the membrane. The elastic wall of the membrane is covered with ion layers, and therefore, a potential difference is created there (the resting potential). In this layer, the concentration and potential change significantly with distance from the membrane due to interactions between the membrane’s ion layers and the electrolyte. Ions sent by other neurons can enter the internal volume through controlled inputs (synapse). Between the membrane and the axon, there is a large amount of uncontrolled ion channels (Axon Initial Segment (AIS)). The conductance of the channels in the membrane’s wall is about 50 times smaller [20, 21] than that in the AIS.
@@
The membrane functions as a Proportional-Integral-Derivative (PID) control circuit. Under a small external influence, a neuron remains in its resting state and regulates itself through non-controlled ion channels in its membrane. If the external influence exceeds a threshold, the controlled channels suddenly release a large amount of ions into the internal space, putting the neuron into a transient state. The membrane potential changes suddenly (within a few dozen nanoseconds), and, due to the mutual repulsion of ions, the pressure acting on the membrane wall and the ion concentration in the layer near the membrane also increase significantly.
@@
Since the same physical action generates the voltage gradient and the concentration gradient (and, consequently, the pressure), they are inseparable and proportional to each other. Generating a voltage change (clamping, direct current, magnetic pulse) or a pressure change (ultrasound, mechanical, shock waves) on the membrane causes the other to change. Inputting ions (synaptic input, followed by rush-in) generates both changes. The control circuit tries to restore the resting state, during which the mentioned changes decrease proportionally together. The transient state of the neuron can, in principle, be described by any of the mentioned entities, but, as a practical matter, the effects of changes in the other entities must also be taken into account: no single discipline can describe the changes. The elastic membrane implements an almost critically damped vibration, but the "vibrating mass" and the "spring force" are constantly changing. In an electrical resonant circuit, the capacitor’s voltage and capacitance change due to the finite amount of charge stored on it.
@@
The pressure wave caused by a force shock and the voltage wave caused by a voltage shock describe the same effect. In both cases, the other effect must be taken into account to some extent: an infinitesimal movement of the ion also involves changes in both pressure and voltage. Changes in pressure and charge alter the membrane’s thickness (and therefore its capacity), and both voltage and pressure can cause a structural change (a phase change, "melting") in the lipid chains that make up the membrane. In the case of the pressure wave, the speed is naturally finite (mass must be moved). For the electrical wave, there are laws regarding only instantaneous interaction; so, they must be adapted to slow ion currents based on physical assumptions. It is simpler to describe the generation of the action potential as an electrical oscillation (although slow currents’ propagation must be accounted for), and its propagation as a pressure wave (although this only represents well the wave front: the AP lasts as long as there is non-equilibrium charge in the neuron: the repulsion between the ions pushes the ions out of the closed space, so the intensity of the wave continuously decays). Hyperpolarization is described in the electrical view as a capacitive current in the sequential oscillator circuit, in the thermodynamic view as a suction effect occurring in the opposite phase of the membrane. In both cases, however, it must be taken into account that charge and mass are inseparable in ions, and that their cross-disciplinary effects significantly reduce the applicability of the disciplinary laws.
@@
In simple words, a neuron combines different disciplines; it works like a two-tact biological internal combustion engine, and can be described theoretically as a special Carnot-type thermodynamic engine. The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a large pressure impulse due to electrical repulsion between ions; see Appendix B. The neuron’s volume remains practically unchanged (represents an iso-volume process). The longitudinal pressure wave vibrates the particles (including the dissociated ions). It is similar to a sound wave, with negligible material transport. The created offset potential moves the ions out of the neuron through the only available output channel: the axon.
In the first stroke, the pressure impulse invests energy into compressing the elastic membrane and starting a compression wave (a soliton). In the second stroke, a combined thermoelectrical/mechanical force (rather than some magic battery) pumps out ions through the ion channels in the AIS, which "resists" it, thereby generating a potential wave known as AP; well described by electricity. The force analysis shows that the pressure wave causes a longitudinal vibration, while the changed potential adds a potential-dependent longitudinal force component for moving the ions.

### v4

@@ Abstract
The paper reviews the conflicting disciplinary theories of neuronal operation and introduces their unified, non-disciplinary synthesis that explains all known experimental observations without contradictions. Although the first principles of science are identical in the inanimate and life sciences, the abstractions and approximations that underlie concrete laws differ, so the actual laws depend on the environment. We challenge the universality of protein mechanisms for describing neuronal operation and show that, by using the appropriate approximations, the non-disciplinary cooperation of thermodynamics and electricity perfectly describes processes in living matter. We show that biology works on the boundary of the micro and macro views of physics, and in some cases, those views must be bridged. Most importantly, we explicitly consider the discrete nature of biological currents, that their charge carriers are slow and they repel each other. That feature, overlooked in both mentioned theories, enables us to interpret the electrical and thermodynamic behavior of important biological objects and processes. The speed of biological currents changes by several orders of magnitude during operation. They keep the electrolyte in continuous movement, making some microscopic regions of the volume responsible for the observed action potential. The cross-disciplinary model enables us to derive the correct electrical equations and thermodynamic description of the action potential. In a way demonstrating the need for the non-disciplinary discussion, we derive the thermodynamic Carnot cycle of AP from measuring electrical parameters, calculating its energy consumption and efficiency, interpreting the neuron’s partially reversible operation, and the long-sought solution to the heat emission/absorption issue. We defy the claim that science cannot describe life: the underlying physics constrains biological phenomena, including cognitive ones.
@@ 1 Introduction
The "great journey into the unknown [understanding the brain]" [2] must begin at a much lower level: revisiting the fundamental phenomena, disciplines, laws, interactions, abstractions, omissions, and testing methods of science. Research must build on classical science, but be reinterpreted for living matter. There is no independent ’life science’, there is only science. It is based on the same ’first principles’ but using different abstractions and approximations for living and non-living matters, and having the appropriate relations between them. "There is a clear need for a tighter and more carefully managed integration and realignment of the work" [3]. Without (re)aligning the knowledge elements along the first principles, the "integration between data and knowledge from different disciplines", lacks integration. However, even the review of the Human Brain Project (HBP) summarized that "HBP is not developing with the expected level of integration and the project controls in place are not adequate to achieve this aim." [3]. We arrived at the boundaries of classical science disciplines and are moving now through terra incognita. We really need a new understanding.
@@
Our model is undoubtedly a physical model, in the above sense. We provide a holistic, multi-level view of neuronal operation and describe (at least semi-)quantitatively the actual physical processes; this way, we achieve the promised ’new understanding’ (and the need for rewriting books in neuroscience). Although the classical neuron model aims to be an electrical model, it violates fundamental laws of electricity and conservation laws; furthermore, it cannot interpret phenomena from other disciplines (thermodynamics, mechanics, optics). The thermodynamic model(s), on the other hand, cannot explain the electrical features of the observed processes. We prove that no compromise is needed: the pressure and the voltage measurements observe the same physical process by different disciplinary means. Neuroscience forgot that neural currents consist of ions, which repel each other in a structured confined volume that contains objects with static charge; furthermore, that ions are charged and have mass, which are abstractions from different classical disciplines.
As Feynman R. P [5] wrote, "the separation of fields [of science] …is merely a human convenience, and an unnatural thing. Nature is not interested in our separations, and many of the interesting phenomena bridge the gaps between fields". "When we go to investigate we should not pre-decide what it is we are trying to do except to find out more about it. …The first principle is that you must not fool yourself, and you are the easiest person to fool." Physiology, despite Hodgkin and Huxley (HH)’s and Feynman’s warning, pre-decided that the only authorized science discipline, electricity, must describe neurons’ operation. However, electricity has no place for the mass of its charge carriers. After seven decades, it should be realized, admitted, and fixed.
@@
We agree that "Biological laws are much more difficult to find than physical ones" [6]. Research already provided the detailed description of a correct ‘net electrical’ neuron model [7], the laws of motion for describing biology [8], and now provides a correct non-disciplinary (unified and scientifically complete electrical/thermodynamic) model. The deeper reason for the apparent conflict between the two competing theoretical descriptions is science’s disciplinarity, as we discuss it separately in Appendix A, since it is somewhat philosophical and touches not only the fundamental concepts of physics but also the way of scientific thinking. Shortly, the disciplinary separation proved to be successful for most phenomena of inanimate science. Electrolytes are an exception. Thermodynamics (due to excluding a long-range interaction between its particles), excludes electric charge (and related quantities) from the quantities it studies. Electricity (due to implying ’electron cloud’ as a charge transmission method) similarly excludes the possibility that long-range interaction may exist between its discrete charge carriers; that is, that they may have thermodynamic features based on the Vlasov equations. This way, electricity excludes mass (and related quantities, such as heat) from the quantities it studies. Neither of the two disciplines, alone, can completely describe electrolytes (and so: biological systems): some of the system’s quantities remain outside the scope of the given discipline. The measurements are not complete: some quantities are not measured, so ’miracles’ can occur. The force analysis does not show a driving force, so some "new force or whatnot" [9] must be assumed (in physiology, "protein mechanisms"), in contrast with science’s principles. Our non-disciplinary approach, however, can successfully describe the operation of a neuron without miracles.
@@
The structure of the paper is as follows. Section 2 presents the state-of-the-art of theoretical neuron modeling. A strongly simplified discussion in section 3, without mathematical, physical, and electrical details, presents the unified model. Electricity of ions and its application to the case of neurons is the subject of section 4, which includes the specifics of ion currents, the electrical features and operation of biological objects, and, first of all, the thermodynamics applied to neurons. Section 5 presents the physical mechanisms that, instead of a magic protein mechanism, define and maintain the membrane’s potential in resting and transient states. The importance of the action potential deserved a separate section 6, where the physical processes, their correct "net electrical" model, with the related mathematics, are described. The discussion culminates in section 7, which on the one side derives the thermodynamic parameters describing Action Potential (AP) and their long sought calculation method, on the other provides evidence that only the cooperation of the competing disciplines can succeed: the thermodynamic parameters can be calculated by measuring parameters of electrical processes.
@@ 2 State-of-the-art
Today, two major branches of disciplinary theories compete for being applied to describe the neuronal operation [10, 11, 12, 13, 14, 15]. Those cited references also compare those theories. The 2-decade-old Heimburg-Jackson thermodynamic theory is trying to break in alongside or displace the 7-decade-old all-electric Hodgkin-Huxley theory. The two theories apply one of the scientific disciplines developed for inanimate nature, unchanged, to living nature. There are unexplained observations for each of the two theories, and neither can describe the biological neuron’s functioning without contradictions. There is no chance of the two theories fusing. Although they can describe a wider range of phenomena together, the conditions under which they are applicable exclude their joint use. Moreover, to conceal contradictions and fill gaps, biophysics assumes mechanisms (typically unspecified protein mechanisms) that contradict the fundamental principles of physics, and some of them even the disciplinary laws.
@@
The electrical view essentially stems from the (Nobel Prize-winning) work of Hodgkin and Huxley [16]. They could not make a perfect job, mainly due to the lack of discoveries made several decades later, including ours, about handling the finite speed of the biological ionic currents (and, due to that, the dynamic features), which drastically change their conclusions. In contrast with their empirical description, which delivered mathematically formulated measured observations, our discussion, although it follows essentially the same principles, reinterprets and pinpoints the used fundamental terms, sets up a physical model, and explains the physically underpinned processes. "However, the theory could not explain the physical phenomena such as reversible heat changes, density changes, and geometrical changes observed in the experiments" [13]. Given that our (non-disciplinary) model natively connects charge and mass of ions [8], it can explain the missing phenomena. The thermodynamic view roots in the (possibly Nobel-prize-winning) idea of Heimburg and Jackson [17]. They proposed that the action potential is essentially a sound wave (a soliton). “However, there are several other questions that this has to answer like ion flow involvement in nerve signal propagation as stated by the HH model and also the faster propagation in myelinated nerves than in unmyelinated” [13]. If we consider that the presence of ions in the neuron’s closed volume means an unbalanced force component exerting only on ions (but not on the vibrating neutral molecules) and the ion flow means charge propagation in a membrane tube where the tube’s specific capacity depends on the thickness of the axonal membrane’s myelin sheath (aka the distance across the membrane surfaces) [18], the faster propagation is not mysterious anymore.
@@
The excellent textbook [19] separates the neuron membrane’s states into resting and transient states. It introduces resting and transient ion channels, but does not introduce that the different physical processes in those states need different models. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; we tell the second half in section 4. We refine their statement that "resting ion channels establish and maintain the resting potential" [19], page 126. The ion channels are passive players. The interplay of the lipid condenser and the electrolytes on both sides of the membrane establishes the resting potential, and ionic gradients maintain it. As we show in section 5, two different physical mechanisms operate in the resting and the transient states, so two different models are needed for describing the same components.
A primary reason neurophysiology has been misguided is that it has a static picture in mind of the dynamic life of neurons. It started from the static inanimate anatomic microscopic pictures, continued with the static clamping investigations (where the natural the negative feedback of an external electric circuit eliminates gradients) and underpinning the theory by analyzing frozen (inanimate) samples by electron microscopes. Those investigations revealed a tremendous amount of crucial details, but they obscured the idea that those samples are just snapshots of a permanently changing system.
Neuroscience projected the mechanisms of the static resting state to the dynamic transient state and is attempting to understand the transient state as perturbations [20] to the (entirely different) resting state. Of course, that attempt required more and more ad hoc non-scientific assumptions. We agree that "Understanding and predicting molecular responses in single cells upon chemical, genetic or mechanical perturbations is a core question in biology." [21] However, it is not possible without understanding that the transient state is a state on its own right, with different physical mechanisms, rather than a perturbation of the entirely different resting state. Claims such as "pumps [by using protein mechanisms] …transport ions against their electrical and chemical gradients" [19], page 101, or "the flux of ions through ion channels is passive, requiring no expenditure of metabolic energy by the channels." [19], page 107, are as scientific as in the pre-Newtonian age, there was a notion related to Aristotelian philosophy. This concept gave credence to the “theory” that the celestial spheres and, later on, the planets were pushed or moved by celestial movers, celestial intelligences, or angels against physical forces.
@@
The book [19] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, discussing only membrane permeability, without explaining how the resting potential is created. We demonstrate in section 5 that classical methods of electricity enable us to calculate the membrane’s potential resulting from charge separation and polarization, and how its resting value is created, furthermore, how it is regulated. We make some physically plausible assumptions to provide numerical figures.
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [22]. We add that mainly physical processes at various speeds (instead of molecular classes such as proteins) and ionic gradients (instead of protein mechanisms) control the processes. More precisely, inseparable thermodynamic and electrical processes set up and maintain the potential established by the electrical and thermodynamic backbone defined by the lipid/protein structure of the cell and the ionic solutions it contains. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [23].
@@ 3 Operation in a nutshell
We have Schrödinger’s comments in mind when constructing the physical picture about neurons, that we must be very careful. We must not use the ’ordinary’ laws of physics derived for non-living matter without revisiting them because "the construction is different from anything we have yet tested in the physical laboratory[9]. Appendix A details the reasons. Although the discussion requires deep knowledge in science, in this section, we remain on the surface and provide a holistic picture of the model. The deeper layers are explained in later sections.
The fundamental principles are the same, but we must use different approximations and abstractions for the different disciplines of science and for living matter. (Notice that when defining the subject and laws of mechanics, we used the approximation that electric charge does not play any role; similarly, in electricity, mass plays no role. Neither of those approximations is valid for ions; so those disciplines, alone, cannot be used to describe electrolytes; consequently, to describe processes in biology.) For this reason, it may be more correct to call those laws governing living matter ’non-disciplinary’ rather than ’non-ordinary’ laws. We must also include the discoveries from the past seven decades, among others, that in the transient state, a neuron shall be modeled as a serial (i.e., differentiator-type) electrical oscillator, which has gated current inputs (instead of a parallel oscillator without gating its inputs). The latter is a good model only in the static view of the resting state.
@@
We explicitly add that the charge carriers, the neuron works with, are ions in electrolytes (instead of electrons in solids) and that the signal propagation mechanism is bioelectric (instead of purely an electrical or mechanical wave or a whatnot "protein mechanism"), in many cases even without having free carriers in the respective volume. (We do not consider the ideas that the associated mechanical, optical, etc. changes cause electric phenomena. They reverse the true causality: they do not have a triggering cause; the view inherited from the narrative biology.) Bioelectrical operation, especially that of semipermeable membranes, requires special attention: charge carriers may be "created" inside the biological matter. Omitting the clear physical background results in performing wrong measurements and introducing wrong concepts into biology purposefully.
3.1 Model components
@@
In the (highly simplified) non-disciplinary physical model (the ’abstract physical neuron’, as we call it), a neuron is an elastic insulating sphere with two lipid layers (membrane), a similar output tube (axon), and a slightly compressible, partially ionized electrolyte fluid with different compositions and concentrations inside and outside. They contain ions, simple chemical molecules, and complex biological formations. These components may have gradients and move at continuously changing speeds that can vary by several orders of magnitude during operation. The lipid layer is partially permeable (controlled or uncontrolled) for the components above (channels, mainly ion channels).
@@
The membrane represents a confined space in which ions behave differently from in an infinite space; furthermore, it significantly alters the properties of the few-nanometer-thick layer near the membrane. The elastic wall of the membrane is covered with ion layers, and therefore, a potential difference is created there (the resting potential). In this layer, the concentration and potential change significantly with distance from the membrane due to interactions between the membrane’s ion layers and the electrolyte. Ions sent by other neurons can enter the internal volume through controlled inputs (synapses). Between the membrane and the axon, there is a large amount of uncontrolled ion channels (Axon Initial Segment (AIS)). The conductance of the channels in the membrane’s wall is about 50 to 100 times smaller [24, 25] than that in the AIS.
3.2 Model operation
@@
The membrane functions as a PID control circuit, but it operates on slow ion currents. Under a small external influence, a neuron remains in its resting state and regulates itself through non-controlled ion channels in the wall of its membrane. If the external influence exceeds a threshold, the controlled channels suddenly release a large amount of ions into the internal space, switching the neuron into a transient state. The membrane potential changes suddenly (within a few dozen nanoseconds), and, due to the mutual repulsion of ions, the pressure acting on the membrane wall and the ion concentration in the layer near the membrane also increase significantly.
@@
Since the same physical action generates the voltage gradient and the concentration gradient (and, consequently, the pressure), they are inseparable and proportional to each other. Generating a voltage change (by clamping, direct current, magnetic pulse) or a pressure change (by ultrasound, mechanical, shock waves) on the membrane causes the other quantity to change. Inputting ions (by synaptic input, followed by rush-in) generates both changes. The control circuit tries to restore the resting state, during which the mentioned quantities decrease proportionally. The transient state of the neuron can, in principle, be described by any of the mentioned entities, but, as a practical matter, the effects of changes in the other quantity must also be taken into account: no single discipline, alone, can describe the changes. The elastic membrane implements an almost critically damped vibration, but the "vibrating mass" and the "spring force" are constantly changing. In an electrical resonant circuit, the capacitor’s voltage and capacitance change due to the finite amount of charge stored on it.
The pressure wave caused by a force shock and the voltage wave caused by a voltage shock describe the same effect: the sudden increase of ions in the layer in question. In both cases, the other effect must be taken into account to some extent: an infinitesimal movement of an ion also involves changes in both pressure and voltage. Changes in pressure and charge alter the membrane’s thickness (and therefore its capacity), and they can cause a structural change (a phase change, "melting") in the lipid chains that make up the membrane.
@@
In the case of a pressure wave, the speed is naturally finite (mass must be moved), but the laws must be adapted to the case of delivering charged particles (that repel each other). For the electrical wave, the laws tacitly assumed instantaneous interaction; so, they must be adapted to slow ion currents. It is simpler to describe the generation of the action potential as an electrical oscillation (although slow currents’ propagation must be accounted for), and its propagation as a pressure wave (although this only represents well the wave front: the AP lasts as long as there is non-equilibrium charge in the neuron: the repulsion between the ions pushes the ions out of the closed space, so the intensity of the wave continuously decays). Hyperpolarization is described in the electrical view as a capacitive current in the sequential oscillator circuit, in the thermodynamic view as a suction effect occurring in the opposite phase of the membrane. In both cases, however, it must be taken into account that charge and mass are inseparable in ions, and that their cross-disciplinary effects significantly reduce the applicability of the disciplinary laws.
@@
In simple words, a neuron combines different disciplines. In the thermodynamic view, it works like a two-tact biological internal combustion engine, and can be described theoretically as a special Carnot-type thermodynamic engine. In the electrical view, it operates as a serial $RC$ circuit (that, of course, has a significant capacitive current in the second phase of operation). The synaptic input currents operate the ignition. The influx of $Na^{+}$ ions acts like an explosion, producing a large pressure impulse due to electrical repulsion between ions. The neuron’s volume remains practically unchanged (an iso-volume process) during the ignition phase. The extension phase is practically limited to the layer near the membrane: ions move slowly. The process is similar to the generation of a sound wave, with negligible material transport.
The force analysis shows that the thermodynamic force, generating a pressure wave, causes a longitudinal vibration, while the changed potential adds a potential-dependent longitudinal force component for moving the ions toward the AIS; as long as there are excess ions in the neuron. The resulting force moves the ions with variable speed out of the neuron through the only available output channel: the axon. The process is entirely in line with the principles of science, but the ion, with its mass and charge, combines two disciplinary forces. After summing the two disciplinary forces, one can use Newton’s Laws to describe the ion’s motion. Notice, however, that the electrical charges move with the speed, which is the bargain between the mechanical wave and the electrical interaction. So, the truth is halfway between the disciplinary explanations: the charge moves under the effect of the resultant force that derives from the vibrating mechanical force exerted on the mass and the electrical repulsion force exerted on the charge of the ions.
In the first stroke, the pressure impulse invests energy into suddenly compressing the elastic membrane and starting a compression wave (a soliton); in the electric view, it suddenly charges the condenser. The two waves are synchronized, see Fig. 18. In the second stroke, a strongly varying combined thermoelectrical/mechanical force (rather than some fixed-voltage magic battery) pumps out ions through the ion channels in the AIS, which "resists" it, thereby generating a potential wave known as AP; well described by electricity. Electricity interprets the backward current as capacitive current. Physiology observes hyperpolarization and explains that the outward current in the channel consists of $Na^{+}$ ions, and when the resultant potential changes its direction, the ions change to $K^{+}$; a kind of alchemy. Alternatively, it hypothesizes intelligent ion channels that synchronize their operation to the operation of ion channels at several centimeters distance regarding the type and amount of ions they deliver, without any evidence for the existence of a communication channel or a control mechanism. Another alternative is that the two kinds of positive ions travel against each other, without a driving force and without their charge acting on each other.